# Demystifying Invariant Effectiveness for Securing Smart Contracts

### v1

@@ 9. Conclusion
In this paper, we present the first comprehensive study of the effectiveness of practical invariant guards on preventing DeFi smart contract attacks. Our large-scale experiments on real-world DeFi hacks demonstrate that the inferred invariant guards are very effective in stopping the existing hacks, but some of the invariant guards can be bypassable by experienced attackers. We also found combining multiple invariants can be more effective than individual invariants with a lower false positive rate.
@@ 10. Data Availability
The experimental data of this paper is available on our website (authors of this paper, 2023). The artifact will be open sourced after the paper is accepted.

### v2

@@ 9. Conclusion
In this paper, we present the first comprehensive study of the effectiveness of practical invariant guards on preventing DeFi smart contract attacks. Our large-scale experiments on real-world DeFi hacks demonstrate that the inferred invariant guards are very effective in stopping the existing hacks, but some invariant guards can be bypassed by experienced attackers. We also show combining multiple invariants can be more effective than individual invariants with a lower false positive rate.
@@ 10. Data Availability
All benchmarks, source code, and study statistics are available at our repositories (Chen et al., 2024c, a, d, b). Additional materials of this paper are available in the website (Chen et al., 2024e).
Acknowledgement
This work was supported by MITACS Accelerate program, NSERC Postdoctoral Fellowship, Singapore Ministry of Education Academic Research Fund Tier 1 (RG12/23), and Nanyang Technological University Centre in Computational Technologies for Finance (NTU-CCTF).